How do I find the indicated term of this geometric sequence??
a1 = 5, r = 2, n = 6

Answer by stanbon(75887)

find the indicated term of this geometric sequence??
a1 = 5, r = 2, n = 6
-----
a(6) = a(1)*r^5
----
a(6) = 5*2^5
----
a(6) = 5*32
----
a(6) = 160
=============
Cheers,
Stan H.
=============